Engineering Entrance Exams

The engine of a truck moving a straight road delivers constant power. The distance travelled by the truck in time t is proportional to A. t B. t 2 C. vt D. t^3/2

abhishek singh
12 Years agoGrade

1 Answer

Pratyay Amrit
12 Years ago
P=`Fv =>mav=k(some constant) =>m(dv/dt)v=k =>v=(t^1/2)*c(some other constant) =>dx/dt=(t^1/2)*c =>x=t^3/2*a(some other contant) Thus, answer is (d)

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago